OpenStudy (anonymous):

Is the sequence 3, 12, 36, ... a geometric sequence? Explain.

OpenStudy (anonymous):

yes because it has a pattern

hartnn (hartnn):

A geometric sequence has common ratio. so is 12/3 = 36/12 ??

hartnn (hartnn):

what is 12/3= ?

OpenStudy (anonymous):

4

hartnn (hartnn):

i asked @magen

OpenStudy (unklerhaukus):